OpenStudy (anonymous):

use quadratic formula to solve. x^2-3x=-5

OpenStudy (anonymous):

0=(1)x^2+(-3)x+(5) therefore a=1, b=-3, c=5 \[x=(\pm \sqrt{b^{2}-4ac})/2a\]

OpenStudy (anonymous):

so x=(sqrt(-11))/2 if your answers are supposed to be real (likely unless you know what complex numbers are) then there is no solution

OpenStudy (anonymous):

thats what I got too but wanted to make sure I didnt do something wrong!

OpenStudy (saifoo.khan):

that is not simplifted.

OpenStudy (saifoo.khan):

simplified*

OpenStudy (anonymous):

what isnt simplified?

OpenStudy (saifoo.khan):

that answer.

OpenStudy (anonymous):

I made the equation x^2-3x+5=0. when I put into the quad equation, I came up with negative 11 under the square root sign. I didn't think you could find the square of a negative number

OpenStudy (saifoo.khan):

\[\sqrt{-11}=i \sqrt{11}\]

OpenStudy (anonymous):

now i understand what the i in the explanation of how to answer the question. so is the answer \[-3\pm i \sqrt{-11}/2\]

OpenStudy (saifoo.khan):

\[\frac{3 \pm i \sqrt {11}}{2}\]

OpenStudy (anonymous):

thats what i meant to type! lol

OpenStudy (saifoo.khan):

haha, yes i knew..